OpenStudy (anonymous):

find the slope of the line passing through the point (-1, -3) and having an x-intercept of 2.

OpenStudy (texaschic101):

slope (m) = (y2 - y1) / (x2 - x1) (-1,-3) x1 = -1 and y1 = -3 (2,0) x2 = 2 and y2 = 0 now we sub slope(m) = (0 - (-3) / (2 - (-1) slope(m) = (0 + 3) / (2 + 1) slope(m) = 3/3 = 1 your slope is 1